How long will it take a 250-N box initially at rest reach a speed of 36 km/h when a force of 65 N is applied on it? The floor is frictionless.




O 1.2 S




O not among the choices




O 2.3 s




O 4.5 s




O 3.9 S

1
Expert's answer
2021-12-06T09:44:46-0500

Given:

"W=250\:\rm N"

"F=65\:\rm N"

"\Delta v=36\:\rm km/h=10\: m/s"

"\Delta t-?"


The Newton's second law says

"ma=F"

Hence, the acceleration of the object

"a=\frac{F}{m}=g\frac{F}{W}=\rm 9.8*\frac{65}{250}=2.55\: m/s^2"

The time of movement

"\Delta t=\frac{\Delta v}{a}=\frac{10}{2.55}=3.9\:\rm s"

Answer: "3.9\:\rm s"